OpenStudy (anonymous):

A regular pyramid has a triangular base with a base edge of 6 inches, a height of 10 inches, and a slant height of 10.33 inches... what is the lateral area? what is the surface area?

OpenStudy (anonymous):

got piture of it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

V[P,10,6,10.33] The volume of a pyramid is equal to (1)/(3) times the area of the length (l) times the width (w) times the height (h). V=(1)/(3)*lwh Substitute the values of the radius (l=6), the width (w=10), and the height (h=10.33) into the formula to find the volume of the pyramid. V=(1)/(3)*6*10*10.33 Cancel the common factor of 3 in the denominator of the first term (1)/(3) and the second term 6. V=(1)/(<X>3<x>)*<X>6<x>^(2)*10*10.33 Reduce the expression by removing the common factor of 3 in the denominator of the first term (1)/(3) and the second term 6. V=1*2*10*10.33 Multiply 1 by 2 to get 2. V=2*10*10.33 Multiply 2 by 10 to get 20. V=20*10.33 Multiply 20 by 10.33 to get 206.6. V=206.6

OpenStudy (anonymous):

dont have pic so not sure does it have a,b,c answers

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well thats the volume of pyrmind just wondering if i got numbers right

OpenStudy (anonymous):

in right order

OpenStudy (anonymous):

what would the lateral area be? what would the surface area be?

OpenStudy (anonymous):

TO GET THE AREA WOULD BE Area = ½b × h b = base h = vertical height

OpenStudy (anonymous):

30 WOULD BE THIS ONE IF 6 IS BASE WITH OR LENGHT AND 10 IS THE HEIGHT

OpenStudy (anonymous):

so 30 would be the surface area? or lateral area?

OpenStudy (anonymous):

30 WOULD BE SURFACE AREA

OpenStudy (anonymous):

I AM WORKING ON LATERAL

OpenStudy (anonymous):

lateral

OpenStudy (anonymous):

The lateral area of a right pyramid can be calculated by multiplying half of the perimeter of the base by the slant height. This is summarized by the formula: LA 5 Ps.

OpenStudy (anonymous):

LA=1/2 PS

OpenStudy (anonymous):

1/2*30*10.33

OpenStudy (anonymous):

YEA

OpenStudy (anonymous):

SO MORE LESS SHOUDL BE 15

OpenStudy (anonymous):

THATS HALF OF THE PERIMTERE

OpenStudy (anonymous):

so lateral is 15?

OpenStudy (anonymous):

WELL SAYS GOT TO TAKE THE PRYMIND AREA =30 SO 1/2 =15 * 10.33

OpenStudy (anonymous):

so 77.48

OpenStudy (anonymous):

double the lengths of the base edge, height, and slant height. what is the ratio of the surface area of the smaller pyramid to the larger pyramid?